A sound wave with intensity 2 × 10-3 W / m^2 is considered as moderately loud to our ears.
The drum inside our ear has a diameter of 6 mm. Show that the energy transferred to the drum of our ear if you hear the above sound wave for 60 seconds is E = 3.4 × 10-6 J. 
Consider the drum as a round surface.
Answer #1

area of the drum =pi*radius^2

=pi*(0.006/2)^2=2.8274*10^(-5) m^2

then power generated=intensity*area

=2*10^(-3)*2.8274*10^(-5)=5.6548*10^(-8) W

energy transferred=power*time

=5.6548*60*10^(-8)

=3.3929*10^(-6) J

=3.4*10^(-6) J